2x+y =13
3x + 3y =24

Answer by mananth(16949)

2 x + y = 13 .............1
3 x + 3 y = 24 .............2

multiply (1)by -3
Multiply (2) by 1
-6 x -3 y = -39
3 x + 3 y = 24
Add the two equations
-3 x = -15
/ -3
x = 5
plug value of x in (1)
2 x + y = 13
10 + y = 13
y = 13 -10
y = 3
y = 3